A painter leans a 26-foot extension ladder against a building, with the
base 10 feet away from a 24-foot wall. The person using the ladder
feels unsafe and moves the base of the ladder 3 feet closer to the wall.
How much shorter will the ladder need to be to reach the same height
on the wall? Round to the nearest foot, if necessary.
Ⓐ 1 foot
Ⓑ 3 feet
Ⓒ 22 feet
Ⓓ 23 feet

Answer by mananth(16946)

A painter leans a 26-foot extension ladder against a building, with the
base 10 feet away from a 24-foot wall. The person using the ladder
feels unsafe and moves the base of the ladder 3 feet closer to the wall.
How much shorter will the ladder need to be to reach the same height
on the wall? Round to the nearest foot, if necessary.
The ladder is shifted 3ft closer to the wall.
The base of ladder now is 7 ft from the wall
The length of ladder = sqrt(24^2+7^2)= 25
So, the ladder will need to be 26−25=1
foot shorter to reach the same height .
.